OpenStudy (anonymous):

Express answer in exact form. Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius. help me please

OpenStudy (saifoo.khan):

this is the same problem.. stuck from sooo many days.. right?

OpenStudy (saifoo.khan):

do u want the answer in pi-form?

OpenStudy (anonymous):

yeah ineed hep

OpenStudy (anonymous):

Area = 64pi - (128pi^3-96root3)/6

OpenStudy (saifoo.khan):

When the radius of the circle is 8" and chord length is also 8" Then this chord forms an equilateral triangle a the center . The angle subtended by the chord/arc at the center is also 60° or π/3 radians Area of the larger segment = Area of the circle − area of the smaller segment Area of the smaller segment = Area of the sector − area of the equilateral triangle formed by the chord at the center Area of a sector = ½ R² θ where θ is the angle in radians subtended by the arc at the center of the circle =½×8²×π/3=32π/3 inch ² Area of equilateral triangle of side 8 "=½×8×8sin60° =32×√3/2 =16√3 inch² Hence Area of the smaller segment = 32π/3 − 16√3 Area of the circle =πR²=π×8²=64π inch² Hence area of larger segment =Area of the circle − area of the smaller segment =64π−{32π/3 − 16√3}= 160π/3 + 16√3 inch²